OpenStudy (anonymous):

find the real solutions of the equation by graphing. x^3-5x^2-12x=-36 (I dont have a graphing calculator)

OpenStudy (anonymous):

you dont need one just graph it the old fashion way

OpenStudy (anonymous):

I still dont know how to do it or anything

OpenStudy (anonymous):

grab a piece of paper pick some points, plug into the equation you get your answer

OpenStudy (anonymous):

 should look like this

OpenStudy (anonymous):

plug in the numbers where im so confused and ive been trying to solve it with a piece of paper and odnt get it

OpenStudy (anonymous):

plug in numbers for x in the equation x^3-5x^2-12x=-36

OpenStudy (anonymous):

a graph has x and y soooo how does that work

OpenStudy (kira_yamato):

Let y = x^3-5x^2-12x+36 I suggest you find the following (a) Turning points --> Set y' = 0 (differentiate y) (b) Nature of turning points --> Plug in values for y'=0 into y'' (differentiate y') if y'' > 0 (min), if y''<0(max) (c) Axial intercepts --> Plug in x=0 for y-intercepts & y = 0 for x-intercepts Plot the turning points, making sure that you know its nature, and axial intercepts. Join them together

OpenStudy (anonymous):

huh?

OpenStudy (kira_yamato):

y = x^3-5x^2-12x+36 y' = 3x^2 - 10x - 12 y'' = 6x - 10

OpenStudy (kira_yamato):

When x = 0, y = 36 When y = 0, x = -3 or 2 or 6 For turning points, let y' = 0 3x^2 - 10x - 12 = 0 \[x = \frac{-5±\sqrt{61}}{3}\]

OpenStudy (anonymous):

ya pretty sure thats not what you need to do...

OpenStudy (anonymous):

the answer is supposed to be 3 individual numbers

OpenStudy (anonymous):

Im so confused